Tom and his attorney are preparing to settle Tom’s personal injury claim. Tom was in a serious auto accident two years ago and is nearly recovered. But his medical bills totaled $30,000, he lost over $20,000 income, spent $7,000 to repair his car, and has spent $3,000 on court costs such as filing fees, transcript costs, and consultation with an expert. (Tom’s attorney fees will not be considered in this exercise.) Tom’s attorney believes there is a 10% chance he could hit a “home run” with a jury and obtain a verdict of $120,000. However, a more likely scenario is that a jury will award Tom all of his out-of-pocket expenses. A worst-case scenario is that a jury will conclude that Tom is 40% at fault for the accident, thus reducing his total out-of-pocket expenses by 40%. Determine Tom’s BATNA along with a reasonable initial offer.
Answer #1

Tom's total out of pocket costs = 30000+20000+7000+3000= 60,000

Tom's Best scenario = 120,000

Tom's worst scenario =0.6*60,000=36000 as 40 % is tom's fault

BATNA is best alternative to negotiated agreement.

Here, Tom's BATNA is 36000

The most likely scenario is 60,000. Hence the inital offer should be above 36000